OpenStudy (anonymous97):

Which of the graphs below correctly solves for x in the equation x2 − 4x + 3 = x + 3?

OpenStudy (anonymous):

You failed to provide a graph

OpenStudy (anonymous97):

Which of the graphs below correctly solves for x in the equation x2 − 4x + 3 = x + 3? Graph of quadratic opening upward and linear sloping up to the right. They intersect at point 0, 3 and point 5, 8. Graph of quadratic opening upward and linear sloping up to the right. They intersect at point 0, 3 and point 5, 8. Graph of quadratic opening downward and linear sloping up to the right. They intersect at point (0,3). Graph of quadratic opening downward and linear sloping up to the right. They intersect at point negative 5, negative 2 and point 0,3.

OpenStudy (anonymous):

however the answers are $5$ and $0$

OpenStudy (anonymous):

so get all the terms to the left side and set it equal to 0 can you do that?

OpenStudy (anonymous97):

x^2-4x=x

OpenStudy (anonymous):

alright but we want it equal to zero so subtract x from both sides and set it equal to 0

OpenStudy (anonymous97):

x^2-3x+3=0?